OpenStudy (anonymous):

Given the equation y − 3 = one half(x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. walk me through this one plz

OpenStudy (anonymous):

i have y-3+one half (x+6) wrote dowm

OpenStudy (phi):

\[ y -3 = \frac{1}{2} (x+6) \] first add +3 to both sides 2nd: distribute the 1/2 over (x+6) 3rd: simplify can you do that ?

OpenStudy (anonymous):

= one half not plus one half

OpenStudy (phi):

to add +3 to both sides of the equation write +3 on both sides then simplify (the left side) to distribute 1/2 , multiply 1/2 times each term inside the parens

OpenStudy (anonymous):

im confused already

OpenStudy (phi):

\[ y -3 = \frac{1}{2} (x+6) \] can you write +3 on both sides ?

OpenStudy (anonymous):